Trying to figure out:
through (6,-1) with m = 3/4
the guide says the answer is:
y = 3/4x - 11/2
I can't figure out how they came to this answer, thank you, tyson

Answer by Cromlix(4381)

Hi Tyson,
Equation of line => y - b = m(x - a)
=> y -(-1) = 3/4(x - 6)
=> y + 1 = 3/4x - 18/4
=> y = 3/4x - 18/4 - 4/4
=> y = 3/4x - 22/4
=> y = 3/4x = 11/2
Hope this helps mate.
:-)